find 3n^2 if n(n+5)=-4
Answer by Nate(3500)

n(n + 5) = -4
n^2 + 5n + 4 = 0
(n + 1)(n + 4) = 0
n = -1 and n = -4
n = -1:
3n^2 = 3(1) = 3
n = -4:
3n^2 = 3(16) = 48